OpenStudy (anonymous):

Day 1 2 3 4 5 6 Cans 50 60 70 75 55 ? Megan is collecting cans for a food drive. The number of cans she has collected so far is shown in the table. Which number of cans collected on day 6 will give Megan’s data a mean absolute deviation of 8.3? A) 50 B) 56 C) 60 D) 66

OpenStudy (anonymous):

I'm not sure but I think it's 50. Basically, for MAD, if I recall correctly, you have to find the average of your input then find the average of the distance of each entry from the average (hence the "absolute"). Can you try it out?

OpenStudy (anonymous):

I got 6.33... for 50.

OpenStudy (anonymous):

What's the formula for mean absolute deviation?

OpenStudy (anonymous):

Are you sure? The average should be 60 and each of the distances are: 10,0,10,15,5,10. I got those, at least.

OpenStudy (anonymous):

\[\Large \begin{array}{rcl} \frac{1}{6}\sum_i^6x_i &=& \mu \\ \frac{1}{6}\sum_i^6|x_i-\mu| &=& 8.3\\ \end{array} \]

OpenStudy (anonymous):

Its basically the average of the absolute difference between the mean and each number.

OpenStudy (anonymous):

\[ \sum_i^6x_i = 50+60+70+75+55 + x_6 \]

OpenStudy (anonymous):

It's kinda tough isn't it. You need \(x_6\) to find the \(\mu\) and \(\mu\) to find \(x_6\)

OpenStudy (anonymous):

Yes, It is 50; could you solve it algebraically?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

I know it is 50; I just don't know how to solve it.

jimthompson5910 (jim_thompson5910):

let x = # of cans collected on day 6 the mean is M = (50+60+70+75+55+x)/6 M = (310+x)/6

OpenStudy (anonymous):

I know the mean is (310 + x)/6 and I subtract that from each value, but because it is absolute value I'm confused

OpenStudy (anonymous):

How do I know which difference is positive or negative?

jimthompson5910 (jim_thompson5910):

Now subtract this mean from each data value and take the absolute value | (310+x)/6 - 50 | = | (10+x)/6 | | (310+x)/6 - 60 | = | (-50+x)/6 | | (310+x)/6 - 70 | = | (-110+x)/6 | | (310+x)/6 - 75 | = | (-140+x)/6 | | (310+x)/6 - 55 | = | (-20+x)/6 | | (310+x)/6 - x | = | (310-5x)/6 |

OpenStudy (anonymous):

What is the absolute value of | (-20+x)/6 | ?

jimthompson5910 (jim_thompson5910):

unfortunately there are two cases | (-20+x)/6 | = (-20+x)/6 if x >= 20 | (-20+x)/6 | = -(-20+x)/6 if x < 20

OpenStudy (anonymous):

So, if there are two cases, how do I determine which case to use?

jimthompson5910 (jim_thompson5910):

the good news however is that | (10+x)/6 | = (10+x)/6 since x > 0

jimthompson5910 (jim_thompson5910):

the rest will have two cases, which may be a pain so the best way to solve is to use a graphing calculator

OpenStudy (anonymous):

I'm not allowed to use a graphing calculator, only a scientific.

jimthompson5910 (jim_thompson5910):

then you'll have to try each case separately

jimthompson5910 (jim_thompson5910):

or you can plug in each answer choice and you'll get the answer that way

OpenStudy (anonymous):

I'm supposed to be able to solve a question similar to this in under 2 minutes. Which way would be faster?

jimthompson5910 (jim_thompson5910):

i guess since your answer choices are at least 50 or over, you can conclude that | (-50+x)/6 | = (-50+x)/6 and | (-20+x)/6 | = (-20+x)/6

jimthompson5910 (jim_thompson5910):

so that saves you a bit of work

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's up?

OpenStudy (anonymous):

When I tried solving the problem, I got (500-4x)/6 = 8.3, but when I solve for x, I get 112.55 which is way off any of the choices.

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

| (310+x)/6 - 50 | = | (10+x)/6 | | (310+x)/6 - 60 | = | (-50+x)/6 | | (310+x)/6 - 70 | = | (110-x)/6 | (because x > 50) | (310+x)/6 - 75 | = | (-140+x)/6 | (because x>50) | (310+x)/6 - 55 | = | (-20+x)/6 | | (310+x)/6 - x | = | (310-5x)/6 |

OpenStudy (anonymous):

Ignore the absolute value symbols.

jimthompson5910 (jim_thompson5910):

Add them all up, then divide that by 6, set that equal to 8.3 [ (x+10)/6+(-50+x)/6+(110-x)/6+(140-x)/6+(-20+x)/6+(310-5x)/6 ] /6 = 8.3 (x+10)/6+(-50+x)/6+(110-x)/6+(140-x)/6+(-20+x)/6+(310-5x)/6 = 8.3*6 (x+10)/6+(-50+x)/6+(110-x)/6+(140-x)/6+(-20+x)/6+(310-5x)/6 = 49.8 I'll let you finish

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so what do you get when you solve for x

OpenStudy (anonymous):

112, I don't know where I messed up in simple addition.

jimthompson5910 (jim_thompson5910):

(x+10)/6+(-50+x)/6+(110-x)/6+(140-x)/6+(-20+x)/6+(310-5x)/6 = 49.8 (x+10-50+x+110-x+140-x-20+x+310-5x)/6 = 49.8 (-4x + 500)/6 = 49.8 -4x + 500 = 49.8*6 -4x + 500 = 298.8 -4x = 298.8-500 -4x = -201.2 x = -201.2/(-4) x = 50.3

OpenStudy (anonymous):

Wait, didn't you already multiply by 6?

OpenStudy (anonymous):

8.3 * 6 * 6 = 298.8

jimthompson5910 (jim_thompson5910):

well for each term you're dividing by 6 (since the mean has you do that) when you find the MAD, you are dividing the total sum by 6

jimthompson5910 (jim_thompson5910):

so that explains why there are 2 instances where you're multiplying by 6

OpenStudy (anonymous):

You kind of lost me there.

jimthompson5910 (jim_thompson5910):

you saw how we got | (310+x)/6 - 50 | = | (10+x)/6 | | (310+x)/6 - 60 | = | (-50+x)/6 | | (310+x)/6 - 70 | = | (-110+x)/6 | | (310+x)/6 - 75 | = | (-140+x)/6 | | (310+x)/6 - 55 | = | (-20+x)/6 | | (310+x)/6 - x | = | (310-5x)/6 | right?

OpenStudy (anonymous):

Wait, nevermind, I think I get it.

jimthompson5910 (jim_thompson5910):

oh ok great

OpenStudy (anonymous):

Each difference was divided by six and you divide the sum of the differences by 6, so you multiply by 36 on both sides of the equation to get rid of the denominators.

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw